The wavelenth of the electromagnetic wave used for AM broadcasting is 110 m.
calculate the frequency of the wave
1
Expert's answer
2018-04-12T09:51:10-0400
The frequency of the electromagnetic wave
f=c/λ=(3×〖10〗^8)/110=2.727×〖10〗^6 Hz=2.727 MHz